# GenSwarm: Scalable Multi-Robot Code-Policy Generation and Deployment via Language Models

### v1

@@ Discussion
This work introduced GenSwarm, an end-to-end system that automatically generates and deploys code policies for versatile multi-robot tasks. As a significant step toward end-to-end generation, GenSwarm presents a novel paradigm that could potentially disrupt the current development process of multi-robot systems. However, GenSwarm has some limitations that could be addressed in the future. First, this study focussed on decision-making and control. Aspects such as sensing and navigation, which are important for practical applications, have not been incorporated. Developing and integrating onboard sensing into the system would be a valuable direction for future research. Second, although GenSwarm demonstrated high success rates for a variety of tasks, there is still potential to improve the success rate and address more complex multi-robot tasks. We believe that achieving these improvements might be challenging when relying solely on LLMs. Therefore, combining language models with other techniques, such as multi-agent reinforcement learning, could be a promising approach. Compared to language models, reinforcement learning is better suited for generating more sophisticated and optimal policies, making it a valuable complement to GenSwarm.
Method
Ten Swarming Tasks
The ten swarming tasks considered in this work are aggregation, flocking, shaping, encircling, crossing, coverage, exploration, pursuing, bridging, and clustering. The following gives the user instructions and evaluation metrics of each task. Multiple metrics may be used to evaluate a task from different aspects. It must be noted that none of these metrics were used when creating code policies using GenSwarm; they are merely used for the purpose of post-evaluation. A task is regarded as successful when all of its corresponding metrics exceed certain predefined thresholds. In this way, we can automatically calculate the success rate of each task. The termination of a simulation trial is triggered when the execution time exceeds certain values or the task has finished in the sense that, for example, all the robots succeed in reaching their desired positions.
  • Aggregation task
    User instruction:The robots need to aggregate as quickly as possible and avoid
    colliding with each other.
@@
    $d_{\text{maxmin}}=\max_{i}\min_{j\neq i}||\mathbf{p}_{i}-\mathbf{p}_{j}||$
    where $||\mathbf{p}_{i}-\mathbf{p}_{j}||$ is the Euclidean distance between robots $i$ and $j$. The task is regarded as successful if the value of this metric is less than 1.
  • Flocking task
    User instruction:The robots must form a cohesive flock, cooperating with all
    others in the environment. The three main behaviors are cohesion, alignment, and
    separation: cohesion maintains connectivity, alignment ensures synchronized move-
    ment, and separation prevents collisions by keeping robots at least 0.5 meters
    apart.
@@
    $\text{Var}_{\text{spat}}=\sum_{d\in\{x,y\}}\text{Var}(P_{d})$
@@
    $d_{\text{DTW}}=\frac{1}{M}\sum_{i<j}\text{DTW}(\mathbf{T}_{i},\mathbf{T}_{j})$
@@
    $\text{DTW}(\mathbf{T}_{i},\mathbf{T}_{j})=\min_{W\in\Omega}\sum_{(a,b)\in W}d\bigl{(}(x_{i}^{a},y_{i}^{a}),(x_{j}^{b},y_{j}^{b})\bigr{)},$
@@
  • Shaping task
    User instruction:The robots need to form a specific shape, with each robot assigned a unique point on that shape. The task requires each robot to move towards and
    maintain its assigned position on the target shape.
@@
    $d_{\text{proc}}=\min_{\mathbf{Q}}\frac{1}{N}\sum_{i=1}^{N}||\mathbf{p}_{i}-\mathbf{Q}\mathbf{p}_{i,\text{T}}||^{2}$
@@
  • Encircling task
    User instruction:The robots need to surround the target prey by evenly distributing themselves along a circle with a radius of 1, centered on the prey. Each robot is
    assigned a specific angle, and they must adjust their positions in real-time based
    on the prey’s movement to achieve coordinated encirclement.
@@
    $d_{\text{error}}=\frac{1}{N}\sum_{i=1}^{N}\big{|}||\mathbf{p}_{i}-\mathbf{p}_{\text{prey}}||-r_{\text{desired}}\big{|}$
@@
  • Crossing task
    User instruction:Each robot must maintain a distance of at least fifteen centi-
    meters from other robots and obstacles to avoid collisions while moving to the
@@
    $\rho_{\text{reach}}=\frac{N_{\text{reach}}}{N}$
@@
  • Coverage task
    User instruction:Divide the environment into sections equal to the number of robots. Each robot needs to move to the center of its assigned section to achieve full
    coverage of the environment.
@@
    $\rho_{\text{area}}=\frac{A_{\text{occupied}}}{A_{\text{total}}}$
@@
    $\text{Var}_{\text{NND}}=\text{Var}(d_{\text{near}})$
@@
  • Exploration task
    User instruction:The robots need to explore all the unknown areas. You are required to assign an optimal sequence of exploration areas to each robot based on the number of robots and the unexplored regions, and then the robots will gradually explore
    these areas.
@@
    $\rho_{\text{visit}}=\frac{N_{\text{visit}}}{N_{\text{total}}}$
@@
  • Pursuing task
    User instruction:Engage in flocking behavior with all robots on the map, moving
    toward the lead robot. The lead robot’s movement is unpredictable, so maintain
    cohesion by staying connected, ensure alignment by moving in sync, and uphold
    separation by keeping a safe personal space. Additionally, be cautious to avoid
    collisions with any obstacles in the environment.
    Evaluation metric: The pursuing task is evaluated based on two metrics. The task is treated as successful when both metrics exceed their corresponding thresholds.
    1) Average distance to prey, denoted as $d_{\text{avg-prey}}$: It measures the average distance between all robots and the prey’s position. It is defined as
    $d_{\text{avg-prey}}=||\mathbf{p}_{\text{avg}}-\mathbf{p}_{\text{prey}}||$
@@
  • Bridging task
    User instruction:The robots need to evenly form a straight line bridge at the
    position where x is equal to zero within the range of y between minus two and two.
@@
  • Clustering task
    User instruction:Robots with initial positions in the same quadrant need to cluster in the designated area of that corresponding quadrant.
@@
    $r_{\text{achieve}}=\frac{\sum_{q=1}^{4}N_{q,\text{achieved}}}{N_{\text{total}}}$
@@
We designed a modular architecture consisting of seven core modules, each containing multiple classes (Extended Data Fig. 2). These classes have inheritance, association, and composition relationships, which enhance system design by enabling code reuse, modularity, and flexibility. The Core Module defines the interfaces between modules, ensuring that they can seamlessly integrate into the system as long as they follow these interfaces. The Skill Module handles skill library operations, including the representation of skills as a skill graph and the functionality to construct, modify, and extend this graph. The Action Module contains all action nodes that encompass tasks such as analyzing constraints, designing functions, writing code, performing syntax checks, and debugging, all guided by interactions with the LLM. The Environment Module encompasses various simulation environments or real-world scenarios, the Constraint Module handles constraint-related tasks, the File Module manages file storage, and the Feedback Module processes all feedback.
@@ Acknowledgments
The authors would like to thank Jialing Lyu for her help in editing the videos. This work was partially supported by the STI 2030-Major Projects (Grant No. 2022ZD0208800). Roderich Groß acknowledges support by the OpenSwarm project, which has received funding from the European Union’s Horizon Europe Framework Program under Grant Agreement No. 101093046.
@@
Extended Data Fig. 1: Different user instructions for the encircling task and some failure cases. a. Comprehensive user instructions result in a relatively high success rate yet are not guaranteed to succeed. b. Concise user instructions result in a relatively low success rate yet may successfully accomplish the task. c. Overly brief user instructions usually result in a low success rate.
@@
Extended Data Fig. 2: The software class diagram of GenSwarm consists of seven modules. The Core Module defines the interfaces between all modules, ensuring seamless integration. The Skill Module manages the skill library and the construction and modification of skill graphs. The Action Module contains action nodes responsible for tasks. The Environment Module supports various simulation and real-world environments, whereas the Constraint Module handles constraint-related tasks. The File Module manages file storage, and the Feedback Module processes all feedback. These modules interact through standardized interfaces defined by the Core Module, ensuring flexibility and consistency across the system.

### v2

@@ Discussion
This work introduced GenSwarm, an end-to-end system that automatically generates and deploys code policies for versatile multi-robot tasks. As a significant step toward end-to-end generation, GenSwarm presents a novel paradigm that could potentially disrupt the current development process of multi-robot systems. However, GenSwarm has some limitations that could be addressed in the future. First, this study focussed on decision-making and control. Aspects such as sensing and navigation, which are important for practical applications, have not been incorporated. Developing and integrating onboard sensing into the system would be a valuable direction for future research. Second, we focus on the framework’s generality and end-to-end automation in this work, rather than the novelty or optimality of the generated policies or collective behaviors. Generating more sophisticated or optimal policies is an important future research topic, which might be challenging to achieve when relying solely on LLMs. Combining language models with other techniques, such as multi-agent reinforcement learning, could be a promising approach. Compared to language models, reinforcement learning is better suited for generating more sophisticated policies, making it a valuable complement to GenSwarm. Third, GenSwarm generates policies from scratch rather than re-using existing ones. This design choice was made to achieve zero-shot capability. Nevertheless, re-using a behavioral repertoire of previously generated solutions is a valuable direction for future research.
Methods
Ten Multi-Robot Tasks
The ten multi-robot tasks considered in this work are aggregation, flocking, shaping, encircling, crossing, coverage, exploration, pursuing, bridging, and clustering. The following gives the user instructions and evaluation metrics of each task. Multiple metrics may be used to evaluate a task from different aspects. It should be noted that these metrics are used solely for post-evaluation but not incorporated into the policy generation pipeline. A task is regarded as successful when all of its corresponding metrics exceed certain predefined thresholds. In this way, we can automatically calculate the success rate of each task. The termination of a simulation trial is triggered when the execution time exceeds certain values or the task has finished in the sense that, for example, all the robots succeed in reaching their desired positions.
Aggregation task: User instruction:The robots need to aggregate as quickly as possible and avoid colliding with each other.
@@ Methods
[…]
where $\mathbf{p}_{i}-\mathbf{p}_{j}||$ is the Euclidean distance between robots $i$ and $j$. The task is regarded as successful if the value of this metric is less than 1.
Flocking task: User instruction:The robots must form a cohesive flock, cooperating with all others in the environment. The three main behaviors are cohesion, alignment, and separation: cohesion maintains connectivity, alignment ensures synchronized movement, and separation prevents collisions by keeping robots at least 0.5 meters apart.
@@
$\text{Var}_{\text{spat}}=\sum_{d\in\{x,y\}}\text{Var}(P_{d})$ (2)
@@
$d_{\text{DTW}}=\frac{1}{M}\sum_{i<j}\text{DTW}(\mathbf{T}_{i},\mathbf{T}_{j})$ (3)
@@
$\text{DTW}(\mathbf{T}_{i},\mathbf{T}_{j})=\min_{W\in\Omega}\sum_{(a,b)\in W}d\bigl((x_{i}^{a},y_{i}^{a}),(x_{j}^{b},y_{j}^{b})\bigr),$ (4)
@@
Shaping task: User instruction:The robots need to form a specific shape, with each robot assigned a unique point on that shape. The task requires each robot to move towards and maintain its assigned position on the target shape.
@@
[…]
@@
Encircling task: User instruction:The robots need to surround the target prey by evenly distributing themselves along a circle with a radius of 1, centered on the prey. Each robot is assigned a specific angle, and they must adjust their positions in real-time based on the prey’s movement to achieve coordinated encirclement.
@@
[…]
@@
Crossing task: User instruction:Each robot must maintain a distance of at least fifteen centimeters from other robots and obstacles to avoid collisions while moving to the
@@
$\rho_{\text{reach}}=\frac{N_{\text{reach}}}{N}$ (7)
@@
Coverage task: User instruction:Divide the environment into sections equal to the number of robots. Each robot needs to move to the center of its assigned section to achieve full coverage of the environment.
@@
$\rho_{\text{area}}=\frac{A_{\text{occupied}}}{A_{\text{total}}}$ (8)
@@
$\text{Var}_{\text{NND}}=\text{Var}(d_{\text{near}})$ (9)
@@
Exploration task: User instruction:The robots need to explore all the unknown areas.
You are required to assign an optimal sequence of exploration areas to each robot
based on the number of robots and the unexplored regions, and then the robots will
gradually explore these areas.
@@
$\rho_{\text{visit}}=\frac{N_{\text{visit}}}{N_{\text{total}}}$ (10)
@@
Pursuing task: User instruction:Engage in flocking behavior with all robots on the map, moving toward the lead robot. The lead robot’s movement is unpredictable, so maintain cohesion by staying connected, ensure alignment by moving in sync, and uphold separa-
tion by keeping a safe personal space. Additionally, be cautious to avoid collisions with any obstacles in the environment.
Evaluation metrics: The pursuing task is evaluated based on two metrics. The task is treated as successful when both metrics exceed their corresponding thresholds.
1) Average distance to prey, denoted as $d_{\text{avg-prey}}$: It measures the average distance between all robots and the prey’s position. It is defined as:
[…]
@@
Bridging task: User instruction:The robots need to evenly form a straight line bridge at the position where x is equal to zero within the range of y between minus two and two.
@@
Clustering task: User instruction:Robots with initial positions in the same quadrant need to cluster in the designated area of that corresponding quadrant.
@@
$r_{\text{achieve}}=\frac{\sum_{q=1}^{4}N_{q,\text{achieved}}}{N_{\text{total}}}$ (12)
@@
We designed a modular architecture consisting of seven core modules, each containing multiple classes (Supplementary Fig. 7). These classes have inheritance, association, and composition relationships, which enhance system design by enabling code reuse, modularity, and flexibility. The Core Module defines the interfaces between modules, ensuring that they can seamlessly integrate into the system as long as they follow these interfaces. The Skill Module handles skill library operations, including the representation of skills as a skill graph and the functionality to construct, modify, and extend this graph. The Action Module contains all action nodes that encompass tasks such as analyzing constraints, designing functions, writing code, performing syntax checks, and debugging, all guided by interactions with the LLM. The Environment Module encompasses various simulation environments or real-world scenarios, the Constraint Module handles constraint-related tasks, the File Module manages file storage, and the Feedback Module processes all feedback.
@@
Code availability
The code of the proposed GenSwarm system is available online: https://github.com/WindyLab/GenSwarm.
@@ Acknowledgments
The authors would like to thank Jialing Lyu for her help in editing the videos. This work was partially supported by the STI 2030-Major Projects (Grant No. 2022ZD0208800) and National Natural Science Foundation of China (Grant No. 62473320, 62473017). Roderich Groß acknowledges support by the OpenSwarm project, which has received funding from the European Union’s Horizon Europe Framework Program under Grant Agreement No. 101093046 and by Robotics Institute Germany (BMBF Grant No. 16ME1001).
@@
Supplementary Information
This file includes Supplementary Figures 1–8 and two Supplementary Movies demonstrating the experimental results.
Supplementary Figures
Supplementary Fig. 1: Performance evaluation for the encircling task subject to sensor noises. Encircling task performance under different noise levels: as noise increases from $\sigma_{0}=0$ to $1.0$, tracking error generally increases.
Supplementary Fig. 2: Performance comparison between different methods. The figure compares our method (GenSwarm) against three baselines (MetaGPT, CaP, LLM2Swarm) and fine-tuned state-of-the-art (SOTA) expert controllers on six tasks over one hundred trials each. All eight metrics are normalized for a lower-is-better evaluation. It can be seen that GenSwarm achieves the best results among the LLM-based methods, and its best-performing policies are competitive with the SOTA controllers.
Supplementary Fig. 3: Success rate comparison between different LLMs. The comparison of four LLM models (GPT-4o, o1-mini, DeepSeek-V3, and Claude-3.7-Sonnet) on multi‑robot task success rates across six representative tasks (shaping, encircling, covering, aggregation, flocking, and crossing). For each combination of model and task, one hundred independent trials were conducted, from user instruction input to policy generation and execution. The average success rates across all tasks for the four models were 71%, 80%, 83%, and 84%, respectively. Although there are minor performance differences, all models perform reasonably robust across the tasks.
@@
Supplementary Fig. 4: Different user instructions for the encircling task and some failure cases. a. Comprehensive user instructions result in a relatively high success rate yet are not guaranteed to succeed. b. Concise user instructions result in a relatively low success rate yet may successfully accomplish the task. c. Overly brief user instructions usually result in a low success rate.
Supplementary Fig. 5: Examples of seven types of prompts for an encircling task. We designed seven prompt types, categorized from unstructured to highly structured. In the figure, green text represents task requirements (objectives and constraints), while orange text represents policy instructions. a. Plain-Compound (Cohesive), which integrates both the task objective and policy into a linguistically coherent paragraph. b. Plain-Compound, which concatenates the objective from Plain-Objective and the policy from Plain-Policy without further linguistic integration. c. Plain-Objective, which provides only the task objective without a policy. d. Plain-Policy, which provides only the policy without the task objective. e. Plain-Narrative, which uses natural, narrative language to describe the task. f. Structured-Objective, which restructures the task requirements into a “description-goal-constraint” format, with no policy. g. Structured-Policy, which explicitly adds policy information within a structured format.
Supplementary Fig. 6: Comparison between different prompt types. The comparison of seven representative prompt types on multi-robot task success rates across three LLM models (GPT-4o, o1-mini, and DeepSeek-V3) and three representative tasks (covering, encircling, shaping). For each combination of prompt type, model, and task, fifty independent trials were conducted, from user instruction input to policy generation and execution. It is observed that prompt types that contained policy details—Plain-Compound(Cohesive) (78%), Plain-Compound (74%), Plain-Policy (74%), and Structured-Policy (74%)—yielded higher success rates. Conversely, prompts lacking this information, such as Plain-Objective (56%), Plain-Narrative (57%), and Structured-Objective (57%), resulted in lower success rates.
@@
Supplementary Fig. 7: The software class diagram of GenSwarm consists of seven modules. The Core Module defines the interfaces between all modules, ensuring seamless integration. The Skill Module manages the skill library and the construction and modification of skill graphs. The Action Module contains action nodes responsible for tasks. The Environment Module supports various simulation and real-world environments, whereas the Constraint Module handles constraint-related tasks. The File Module manages file storage, and the Feedback Module processes all feedback. These modules interact through standardized interfaces defined by the Core Module, ensuring flexibility and consistency across the system.
Supplementary Fig. 8: Screenshot of an experimental video (Movie 1). View 1 shows a screen recording of the computer’s terminal, where real-time logs are displayed and a user provides natural language inputs. View 2 shows an overview of the physical setup, including the user station and the multi-robot platform. View 3 shows a top-down view of the robots executing the task.
Supplementary Movies
Supplementary Movie 1. One-take end-to-end demonstration of GenSwarm performing the encircling task.
Supplementary Movie 2. One-take end-to-end demonstration of GenSwarm performing the flocking task.